show all work on paper
A. what the variable represents.
b. an equation that represents the equation.
c. the solution of the equation.
d.a complete sentence that answers the question in the problem.
1. the difference of two numbers is 6, and their product is 160. what are the two numbers?
thankyou, brandon stone email is stonewoodfarms87@aol.com i have come up woth 6 and 10 but dont know if its correct nor do i know how i would put it into an equation... am i correct?

Answer by flame8855(424)

let x and y be the two numbers
x-y= 6
xy=160
x=6+y
y(6+y)=160
y^2 + 6y -160 =0
y =10
xy=160
y =16
the two numbers are 16 and 10